# Growth of Light-Seed Black Holes in Gas-rich Galaxies at High Redshift

### v1

@@ 5 Conclusions
In this work, we investigated whether the conditions necessary for light seed growth occur within gas-rich galaxies at high redshift. Using simulations performed with the Arepo code (Springel, 2010; Pakmor et al., 2016), we modeled the creation and accretion behaviour of sink particles, which represent stars during their lifetimes and later evolve into BHs.
The initial galaxy, with a mass of approximately $10^{8}\ \text{M}_{\odot}$, was generated using GalIC (Yurin & Springel, 2014), allowing for a stable and self-consistent initial condition. Moreover, the galaxy setup we employ models the chaotic, gas rich and relatively compact galaxies expected at high-z. Two realisations of this galaxy were explored: one where SNe feedback was disabled, and another where SNe feedback was enabled. These two configurations allowed us to assess the impact of SNe feedback on the environments and subsequent accretion that could support light seed formation and growth. The following are our conclusions:
@@
  • In the realisation with no-feedback, we formed $2004$ MBH particles among which $1520$ were able to accrete more than $1$ $\ \text{M}_{\odot}$. $33$ MBH particles grew larger than $10^{5}$ $\ \text{M}_{\odot}$. The fastest growing MBH particles grew larger than $10^{5}$ $\ \text{M}_{\odot}$ within just $\sim 10^{4}$ years. Such rapid accretion is observed because these particles do not decouple from the gas clump they were formed from. In the absence of feedback, they rapidly accrete the entire gas clump. These MBH particles have initial Eddington factors between $10^{2}-10^{4}$, and the BHs grow to $\sim 10^{5}$.
  • The realisation with feedback formed $2082$ MBH particles (around a 100 PopIII particles underwent PISN and were deleted) and $184$ out of them accreted more than $1\ \text{M}_{\odot}$. The addition of feedback decreased the number of growing MBH particles by an order of magnitude. However, $16$ MBH particles were still able to grow beyond $10^{5}$ $\ \text{M}_{\odot}$. Some of these particles formed through the direct collapse channel, so they experienced no feedback. For the remaining, we find that the SNe feedback was inefficient in pushing back the surrounding gas, causing it to recollapse on the MBH particles. This expansion and recollapse delays the growth to $10^{5}\ \text{M}_{\odot}$ but does not stop it.
@@
In summary, we find that even in the presence of SNe feedback, light seed growth is possible for approximately 1% of BHs in dense, gas rich, environments at high redshift. The total halo mass of our realisation is approximately $10^{8}$ $\rm{M_{\odot}}~{}$and these galaxies are typical of proto-galaxies expected at high-z. Given the right external conditions such galaxies are likely ideal incubators for growing the first MBHs starting from a light seed. In a forthcoming study we will place these results in a cosmological context and compare heavy seed number densities directly against light seed number densities with the goal of determining the breakdown of which environments will fulfill the requirements of light seed growth vs. heavy seed formation and growth.

### v2

@@ 6 Conclusions
In this work, we investigated whether the conditions necessary for light seed growth occur within gas-rich galaxies at high redshift. Using simulations performed with the Arepo code (Springel, 2010; Pakmor et al., 2016), we modelled the creation and accretion behaviour of sink particles, which represent stars during their lifetimes and later evolve into BHs.
The initial galaxy, with a mass of approximately $10^{8}\ \text{M}_{\odot}$, was generated using GalIC (Yurin & Springel, 2014), allowing for a stable and self-consistent initial condition. Moreover, the galaxy setup we employ models the chaotic, gas rich and relatively compact galaxies expected at high-z. Two realisations of this galaxy were explored: one where SNe feedback was disabled, and another where SNe feedback was enabled. These two configurations allowed us to assess the impact of SNe feedback on the environments and subsequent accretion that could support light seed formation and growth. In this work, we have neglected radiative feedback from PopIII stars. We also employed a stellar lifetime shortening to properly capture the BH phase of a sink particle. Both of these caveats may alter our results and reduce the number of growing BHs or their final masses, we will investigate the impact of both PopIII radiative feedback and BH feedback in future works. The following are our conclusions.
@@
  • In the realisation with no-feedback, we formed $2004$ MBH particles among which $1520$ were able to accrete more than $1$ $\ \text{M}_{\odot}$. $33$ MBH particles grew larger than $10^{5}$ $\ \text{M}_{\odot}$. The fastest growing MBH particles grew larger than $10^{5}$ $\ \text{M}_{\odot}$ within just $\sim 10^{4}$ yr. Such rapid accretion is observed because these particles do not decouple from the gas clump they were formed from. In the absence of feedback, they rapidly accrete the entire gas clump. These MBH particles have initial Eddington factors between $10^{2}-10^{4}$, and the BHs grow to $\sim 10^{5}$.
  • The realisation with feedback formed $2082$ MBH particles (about 100 PopIII particles underwent PISN and were deleted) and $184$ of them accreted more than $1\ \text{M}_{\odot}$. The addition of feedback decreased the number of growing MBH particles by an order of magnitude. However, $16$ MBH particles were still able to grow beyond $10^{5}$ $\ \text{M}_{\odot}$. Some of these particles formed through the direct collapse channel, so they experienced no feedback. For the remainder, we find that the SNe feedback was ineffective in pushing back the surrounding gas, causing it to recollapse on the MBH particles. This expansion and recollapse delay growth to $10^{5}\ \text{M}_{\odot}$ but does not stop it.
@@
In summary, we find that even in the presence of SNe feedback, light seed growth is possible for approximately 1% of BHs in dense, gas rich environments at high redshift. The total halo mass of our realisation is approximately $10^{8}$ $\rm{M_{\odot}}~{}$and these galaxies are typical of proto-galaxies expected at high-z. Given the right external conditions, such galaxies are likely ideal incubators for growing the first MBHs starting from a light seed. In a forthcoming study we will place these results in a cosmological context and compare heavy seed number densities directly against light seed number densities with the goal of determining the breakdown of which environments will fulfil the requirements of light seed growth vs. heavy seed formation and growth.